OpenStudy (anonymous):

What is the standard form of (1 + 7i)(3 − i)? a.(-10 − 10i) b.(-4 − 10i) c.(10 + 20i) d.(20 + 10i) @marie1738

OpenStudy (abdullahm):

We need to use the FOIL method.  Also, remember that i * i = i^2 = -1